How can I rewrite the equation 4x – 6y = –30 as a function of x.
Answer by funmath(2925)

Rewrite the equation 4x – 6y = –30 as a function of x.
-4x+4x-6y=-4x-30
0-6y=-4x-30
-6y=-4x-30
-6y/-6=-4x/-6-30/-6
y=(2/3)x+5
Now y is a function of x. y depends on the value of x.
Happy Calculating!!!